# DocNet: Semantic Structure in Inductive Bias Detection Models

### v1

@@ 6 Limitations
This project primarily relies on partisan topics and does not look into bias detection along non-partisan spectrums. In addition, as the data was collected based on what Tweets on the same topic were shared, there may be a skew in our dataset to highlight media with characteristics that draw the attention of Twitter users. However, given the size of the dataset, the variation in domains present, and the consistency in trends with the manually labeled BASIL dataset, we believe our data to be sufficiently generalizable. Without further data validation, our data should only be used for research purposes. An incorrect bias detection model may lead to misplaced mistrust or over confidence in media. This project also only analyzes articles in the English language. While we believe our methodology is language agnostic, this hypothesis has not yet been validated on news articles from other languages.
Disclaimer
The views expressed in this paper are those of the authors and do not reflect the official policy or position of the United States Military Academy, the United States Army, the Department of Defense, or the United States Government.
7 Ethics Statement
All articles collected for this study were done so under the provisions of Section 107 of the U.S. Copyright Act and ensured that our collection action fell under the fair use category. The Tweets were collected in accordance with Twitter’s terms of service at the time of collection. The code and news articles are made available at https://github.com/nlpresearchanon/DocNet under the MIT and CC0 1.0 Universal licenses, respectively.

### v2

@@ 6 Limitations
This project relies on partisan topics and does not look into bias detection of non-partisan capacities. In addition, as the data was collected based on what Tweets on the same topic were shared, there may be a skew in our dataset to highlight media with characteristics that draw the attention of Twitter users. However, given the size of the dataset, the variation in domains present, and the consistency in trends with the manually labeled BASIL dataset, we believe our data to be sufficiently generalizable. Without further data validation and given copyright restrictions, our data should only be used for research purposes. An incorrect bias detection model may lead to misplaced mistrust or over confidence in media. Additionally, this project also only analyzes articles in the English language. While we believe our methodology is language agnostic, this hypothesis has not yet been validated on news articles from other languages.